OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#11568 — Optical loop Marseille
Incident Report for Web Cloud
Resolved
We have a fiber cut between SLO13 and PoP OVH. There is no impact on traffic. The NRA of SLO13 and SMT13 are no longer secure, optical security takes over.

Update(s):

Date: 2014-09-15 02:24:51 UTC
Fixed.

Date: 2014-09-15 02:22:44 UTC
The problem is being handled by the provider via a team that is currently intervening on the fiber.
Posted Sep 15, 2014 - 02:21 UTC